2. In a hypothetical economy, farmer grows wheat, which he sells to a miller for Rs 40,000. The miller

turns the wheat into flour, which he sells to a baker for Rs 70,000. The baker turns the flour into

bread and sells to consumers for Rs 1,20,000. Consumers eat the bread.

a) Using Value added approach, compute GDP by mentioning value added at each activity level.

b) Who contributes more to GDP, farmer, miller or baker and how much?

c) Does this example suggest another way of computing GDP, other than Value added approach? If

yes, suggest the name of the approach, compute GDP and compare the result which you found

with the value of GDP computed by Value added approach.


1
Expert's answer
2021-12-09T09:43:51-0500

(a)

Farmer to Miller: Rs 40000

Miller to Baker: Rs 70000

Baker to consumer: Rs 120000

Value added by miller "=70000-40000=30000"

Value added by baker:"=120000-70000=50000"

GDP is the final value sold to the consumer:"=Rs 120000"

"=40000+30000+50000"

"=Rs 120000"


(b)

The baker contributes more to GDP, value added by baker is Rs 50000.


(c)

This example does not suggest any other way of computing GDP apart from value added approach.

The other alternatives for calculation of GDP are the income approach and the expenditure approach methods which do not apply in this case.